What is the remainder? x4 - 3x2 + 5x is divided by 2x - 1


1
Expert's answer
2021-12-13T07:23:55-0500

"\frac{x^4-3x^2+5x}{2x-1}=\frac{x^3}{2}+\frac{x^2}{4}-\frac{11x}{8}+\frac{29}{16}+\frac{\frac{29}{16}}{2x-1}."

Remainder: "\frac{29}{16}."